OpenStudy (anonymous):

80% of what number is 16?

OpenStudy (anonymous):

80/100 * n = 16 n =

OpenStudy (anonymous):

n = (16 * 100)/ 80 n = 20

OpenStudy (anonymous):

\[0.8 \times x = 16\]

OpenStudy (anonymous):

\[x = 16\div 0.8\]

OpenStudy (anonymous):

\[x = 20\]

OpenStudy (anonymous):

i do not know

OpenStudy (anonymous):

LOL THANK YOU MRS~HATE~BIEBER

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

OK I GET 0.8 AND THEN I DIVIDE IT BY WAT?

OpenStudy (anonymous):

SO IS THIS 10? 40% of what number is 32?

OpenStudy (anonymous):

youv'e posted that like 19 times

OpenStudy (anonymous):

Use @johanprinsloo's equation: .4 * x = 32 x = 32 / .4 x = 80